The reaction below, carried out at 272.67 K and 1.00 bar, carried out at constant pressure was found to release 329 kJ of heat and consume 21.8 g of oxygen.
Calculate the internal energy change associated with this reaction.
2H2(g) + O2(g) → 2H2O(g)
a. 331 kJ
b. 329 kJ
c. -331 kJ
d. -327 kJ
e. -329 kJ
Answer #1

According to 1st law of thermodynamics ;\DeltaU = q + W

where \DeltaU is internal energy change

q is the heat release or gain to system

W is the work done

w=p*V =nRT as per ideal gas equation

n= no of moles = 21.8/32= 0.68125

W= 0.68125mole*272.67K* 8.314JK-1mol-1= 1544.37 J= 1.544kJ

\DeltaU = -329kJ( negative sign. for heat releasing) + 1.544kJ = -327.455kJ = -327kJ